OpenStudy (lucaz):

if cos(x)=1/3 what is sin(x)?

OpenStudy (lucaz):

how to find it?

terenzreignz (terenzreignz):

A triangle should help you...

OpenStudy (lucaz):

how come?

terenzreignz (terenzreignz):

 Here's a right triangle, with the angle x.

terenzreignz (terenzreignz):

Now, you're given cos(x) = 1/3 which two sides does the cosine-function relate?

OpenStudy (lucaz):

adj/hyp

terenzreignz (terenzreignz):

good. so, the adjacent side measures...

OpenStudy (lucaz):

1/3

terenzreignz (terenzreignz):

No...well maybe, but it's adj/hyp and 1/3 so which is adj?

OpenStudy (lucaz):

oh, 1

terenzreignz (terenzreignz):

good, and hyp is 3. Let's mark those out Can you find the length of the missing side? Use the Pythagorean theorem.

OpenStudy (lucaz):

lol, I see it, thanks!

terenzreignz (terenzreignz):

Really? 'Coz I don't :> LOL Post your answer, but remember, that missing side is NOT the answer yet, you're asked for sin(x) ;)

OpenStudy (lucaz):

no, it's ok, I wasn't recognizing the cos as 1/3, I can do rest of my problem now

terenzreignz (terenzreignz):

Well, if you're sure, then, great ^_^ Though, do post your answer, if you're willing :>

OpenStudy (lucaz):

3^2=1^2+?^2.....

OpenStudy (lucaz):

I was stuck with the 1/3, thank you

terenzreignz (terenzreignz):

oh okay. Well, you seem to know what you're doing. ^_^